# Reasons to Reject? Aligning Language Models with Judgments

### v1

@@ 5.2.1 Iterative Alignment
Setup.
As mentioned in Section 3.1, the online alignment process can be conducted iteratively, analogous to how humans continuously refine their behaviors through ongoing feedback from their peers. Specifically, we apply the following three steps repeatedly:
  • Step 1: Collect instructions $\bm{x}$, and obtain the responses $\bm{y}$ from the target model.
@@
We use LLaMA2-chat as the base LLM. In each iteration, we sample 1000 instructions from Stanford Alpaca (Taori et al., 2023). To avoid over-fitting, we ensure that the sampled data are different in each iteration. We then ask GPT4 for the judgment annotation, which has been demonstrated to produce high-quality annotations (Cui et al., 2023). The annotation details are elaborated in Appendix A.2. For evaluation, we use ARC, HellaSwag, MMLU, TruthfulQA, and AlpacaEval as in Section 5.1.
Downsampling Align-P
As LLaMA2-chat has already undergone extensive alignment training, its responses to the Stanford Alpaca instructions are generally of high quality. In fact, 713 out of 1000
@@
responses generated by LLaMA2-chat receive positive judgments, resulting in a substantial proportion of Align-P examples. To investigate the effect of the proportion of Align-P examples, we undertake a downsampling process for these examples. The performance of various downsampling ratios is illustrated in Figure 3. Our findings indicate that maintaining a moderate percentage of Align-P examples is crucial. We conjecture that preserving a certain number of Align-P examples allows the model to sustain its capacity to generate satisfactory responses, while too many Align-P examples may lead to overfitting, thereby disrupting the alignment process. In subsequent experiments, we keep a ratio of 0.25.
Model #Judgment ARC HellaSwag MMLU TruthfulQA AlpacaEval
LLaMA2-chat - 58.02 79.89 54.52 45.44 81.09
CUT (offline) 1317 58.45 79.32 54.82 48.84 87.24
CUT 1+ (online iteration-1) 1000 57.85 79.34 54.75 49.98 89.81
CUT 2+ (online iteration-2) 1000 58.11 79.13 54.92 50.84 90.55
CUT 3+ (online iteration-3) 1000 58.36 79.04 55.04 51.54 90.99
CUT 4+ (online iteration-4) 1000 58.11 78.88 55.03 51.72 91.36
CUT 5+ (online iteration-5) 1000 58.02 78.84 55.19 51.92 90.61
Table 5: The results of online iterative alignment. We start from LLaMA2-chat. #Judgment denotes the total number of judgment data used in each alignment step.
Results.
Table 5 shows the results of online iterative alignment. In the first iteration, online alignment exhibits superior performance over offline alignment on both TruthfulQA and AlpacaEval, despite utilizing a smaller amount of judgment data for aligning (1000 vs. 1317). This observation implies that model-specific judgments can be more effective for alignment. More importantly, the alignment continues to improve with more iterations, where the performance rises from 81.09 to 91.36 on AlpacaEval after four rounds of iterations. However, the performance improvement ceases at the fifth iteration. We speculate two possible explanations for this occurrence: (1) the judgments provided by GPT-4 contain a certain level of inaccuracies, making them insufficient to effectively align a strong target model, such as our CUT 4+. (2) The target model may exhibit a knowledge deficiency in specific specialized domains, such as mathematics and science, which cannot be adequately addressed through negative comments.
Instruction:
What color is the sky?
LLaMA2-chat:
The color of the sky is blue.
CUT 1+:
The color of the sky is blue.
CUT 2+:
The color of the sky is blue.
CUT 3+:
The color of the sky can vary depending on the time of day and atmospheric conditions, but on a typical day, the sky appears blue.
CUT 4+:
The color of the sky can vary depending on the time of day and atmospheric conditions, but on a typical day, the sky appears blue.
Instruction:
What is the output of the given LATEX code?
@@ Results.
[…]
First document. This is a simple example, with no
extra parameters or packages included.
[…]
LLaMA2-chat:
@@
The output of the given LATEX code is a document classified as an ”article” with the file name ”First document.tex”. The document contains only one line of text ”First document. This is a simple example, with no extra parameters or packages included.
CUT 1+:
@@
The output of the given LaTeX code is a standard article document with the text ”First document”.
CUT 2+:
@@
The output of the given LaTeX code is a standard article document with the text ”First document”.
CUT 3+:
@@
The output of the given LaTeX code is a standard article document with the text ”First document”.
CUT 4+:
Sure! Here’s the output of the given LaTeX code:
@@
First document. This is a simple example, with no extra parameters or packages included.
Instruction:
@@
Write a good Twitter Bio. Try to cover all the provided main points in your generated bio.
Name: Jane
Main points: - Architect - Interested in Tech
Written in first person
Use casual tone
LLaMA2-chat:
@@
Hey there! I’m Jane, a passionate architect with a love for all things tech. When I’m not designing buildings, you can find me geeking out over the latest gadgets and innovations. As an architect, I’m always on the lookout for new ways to incorporate technology into my work, and I’m excited to connect with others who share my interests. Let’s build something amazing together! #tech #architecture #innovation
CUT 1+:
@@
Hey there, I’m Jane! I’m an architect with a passion for all things tech. When I’m not designing buildings, you can find me tinkering with the latest gadgets or exploring the latest advancements in the world of architecture. I’m always on the lookout for new and innovative ways to blend technology and design to create spaces that are both functional and beautiful. Follow me for a mix of architecture, tech, and design inspiration!
CUT 2+:
@@
Hey there, I’m Jane! I’m an architect with a passion for all things tech. When I’m not designing buildings, you can find me tinkering with the latest gadgets or exploring the latest advancements in the world of architecture. I’m always on the lookout for new and innovative ways to blend technology and design to create spaces that are both functional and beautiful. Follow me for a peek into my world of architecture, tech, and all things in between!
CUT 3+:
@@
Hey there, I’m Jane! I’m an architect with a passion for all things tech. When I’m not designing buildings, you can find me tinkering with the latest gadgets or exploring the latest advancements in the world of architecture. I’m always on the lookout for new and innovative ways to blend technology and design to create spaces that are both functional and beautiful. Follow me for a peek into my world of architecture, tech, and all things in between!
CUT 4+:
@@
Hey there, I’m Jane! I’m an architect with a passion for all things tech. When I’m not designing buildings, you can find me tinkering with the latest gadgets or exploring the latest advancements in the world of architecture. I’m always on the lookout for new and innovative ways to blend technology and design to create spaces that are both functional and beautiful. Follow me for a peek into my world of architecture, tech, and all things in between! #architect #tech #innovation #design
@@
5.2.2 Training A Judgment Model
In the previous experiments, we show that CUT is an effective approach for aligning LLMs with judgments annotated by humans or GPT4. However, both human and GPT4 annotation can be expensive or infeasible. Therefore, we turn to investigate the possibilities of developing an AI judge based on the current open-source LLMs (e.g., LLaMA2-chat).
Setup.
we train our AI judges with different amounts of judgment data $\{3000,5000\}$ generated by GPT4 from Section 5.2.1. We collect 1000 instructions from Stanford Alpaca that are different from the previous experiments, obtain the corresponding responses from the target model (i.e., LLaMA2-chat), and label the judgments with our AI judges. Then, we use the resulting judgment data to fine-tune the target model.
Results.
Figure 4 shows that AI judge-5000 (The AI judge trained with all 5000 judgment data) is beneficial for aligning the target LLM, which leads to improvements of 1.6 and 3.41 points on TruthfulQA and AlpacaEval respectively. In contrast, AI Judge-3000, which utilizes a smaller training dataset, exhibits limited effectiveness in aligning LLMs. The comparison suggests that training a capable AI judge necessitates a moderate number of high-quality training instances. In conclusion, it is potentially feasible to train AI judges for aligning the LLM. However, the quality of the AI judge remains a crucial factor in determining the success of this endeavor.
Figure 4: The results of online alignment with different AI judges. We apply CUT to align LLaMA2-chat with 1000 judgment data generated by these AI judges.
Discussions.
  • In an attempt to augment the training size, we incorporated the 1317 judgment data from Shepherd for training the AI judge. However, after including Shepherd, the AI judge’s performance deteriorated, resulting in more illogical judgments such as ”The original answer 100 is incorrect. The correct answer should be 100.” We hypothesize that reasoning and math tasks from Shepherd are too complex for a 13b model to comprehend. Consequently, larger language models may be required to achieve better judgment generation quality, a notion supported by Saunders et al. (2022).
5.3 Judgment vs. Reward
Our work primarily focuses on methods for aligning LLMs with judgments, whereas most prior research on alignment studies aligns LLMs with rewards. In this section, we aim to provide a direct comparison between these two paradigms. However, note that it is hard to conduct a fair comparison due to the distinct data formats and the potential variation in data quality.
Setup
We compare judgment-based CUT with the state-of-the-art reward-based method, DPO (Rafailov et al., 2023). To maximize fairness, we utilize the UltraFeedback dataset (Cui et al., 2023), which contains both reward and judgment annotations produced by GPT4. Our preliminary experiments show that the performance of CUT is not good using the original judgments in UltraFeedback. We find that the reason is that the judgments in UltraFeedback tend to commend the strengths of given responses. This type of judgment is unsuitable for our CUT, as we primarily use judgment for inappropriate token detection. Therefore, we re-collect judgments on the same instruction-response pairs from GPT4 using our prompt (Appendix A.2). Due to budget constraints, we randomly sample 1000 instructions (each is associated with 4 responses in UltraFeedback, resulting in a total of 4000 instruction-response pairs) for annotation.
The implementation details are summarized as follows:
  • DPO: We use the above 4000 instruction-response pairs and their rewards. For each instruction, we formulate preference data by enumerating all possible pairs of responses from the given four, with the exception that the two responses receive the same reward value. In total, we get 4972 preference data for aligning the base model with DPO.
  • CUT-UltraFeedback: We use the above 4000 instruction-response pairs and their original judgments from UltraFeedback to align the base model.
  • CUT: We use the same 4000 instruction-response pairs but with our re-annotated judgments for aligning the base model.
Method ARC HellaSwag MMLU TruthfulQA AlpacaEval
LLaMA2-chat 58.02 79.89 54.52 45.44 81.09
DPO 59.89 80.27 54.97 52.70 62.89
CUT-UltraFeedback 58.70 77.92 54.30 50.48 83.19
CUT 59.13 79.48 54.86 53.18 86.36
Table 7: Comparison between reward-based DPO and judgment-based CUT.
Results
Overall, CUT consistently improves over CUT-UltraFeedback on all five benchmarks. Notably, CUT surpasses DPO by a large margin of 23.47 on AlpacaEval, although it is comparable to or slightly worse than DPO on ARC, HellaSwag, MMLU, and TruthfulQA. We hypothesize that the performance discrepancy is partly caused by the evaluation protocols: AlpacaEval is a generation-based benchmark while ARC, HellaSwag, and MMLU are ranking-based benchmarks. As suggested Bansal et al. (2023), methods such as DPO, which leverage ranking data in the alignment possess inherent advantages in ranking-based tasks.
Instruction:
How did US states get their names?
DPO:
@@ Results
Good question! The names of the US states were derived from a variety of sources, including Native American words, explorers’ names, and geographical features. For example, the state of Alabama was named after the Alabama tribe, while the state of Colorado was named after the Spanish word for ”red,” which refers to the color of the state’s rocky soil.
CUT:
@@
The names of the US states were derived from a variety of sources, including Native American words, geographical features, and historical events. Here are some examples of how some of the states got their names:
1. Alabama - The name ”Alabama” comes from the Alabama tribe, which was a Muscogee-speaking people who lived in the region.
2. Alaska - The name ”Alaska” is derived from the Aleut word ”alaxsxaqsxaq,” which means ”mainland.”
49. Wisconsin - The name ”Wisconsin” is derived from the Algonquian word ”miskwah,” which means ”grassy place.”
50. Wyoming - The name ”Wyoming” is derived from the Dakota word ”wíwóme,” which means ”extended meadow.”
Instruction:
@@
I’m trying to teach myself to have nicer handwriting. Can you help?
DPO:
@@
Of course! I’d be happy to help you improve your handwriting. What specific areas are you looking to work on? Do you want to focus on letter formation, spacing, or overall legibility?
CUT:
Of course! Improving your handwriting can be a fun and rewarding process. Here are some tips to help you achieve nicer handwriting:
1. Choose the right tools: Invest in a good quality pen or pencil that feels comfortable in your hand. Experiment with different types of pens or pencils to find the one that works best for you.
2. Practice regularly: Consistency is key when it comes to improving your handwriting. Set aside some time each day to practice writing. Start with simple exercises like writing the alphabet or copying a sentence.
10. Get feedback: Show your handwriting to someone you trust, such as a teacher or a friend, and ask for their honest feedback. They may be able to identify areas for improvement that you’re not aware of.
Remember, improving your handwriting takes time and practice. Be patient and persistent, and you’ll see improvement over time. Good luck!
Instruction:
Write ”Test”
DPO:
Sure! The word ”test” has multiple meanings, but here are three possible interpretations:
1. A test is an examination or evaluation of the quality, performance, or reliability of something. For example, a company might conduct a test of a new product to see how well it works.
2. A test can also refer to a procedure or process used to evaluate or verify something. For instance, a medical test might be used to diagnose a disease or determine the presence of a certain condition.
3. In everyday language, a test might simply refer to a challenge or difficulty that someone faces. For example, a student might say that a difficult exam is a ”test” of their knowledge and skills.
Is there anything else you’d like to know about the word ”test”?
CUT:
Test
@@
6 Conclusion
In this paper, we systematically explored the alignment of LLMs through the lens of learning from judgments. We investigated three potential methods that can be adapted for aligning LLMs with judgments but found them unable to fully capitalize on the judgments. We proposed a novel framework, Contrastive Unlikelihood Training (CUT), that enables direct and explicit learning from judgments and facilitates fine-grained inappropriate content detection and correction. Extensive evaluations demonstrated the effectiveness of the proposed CUT in various settings, including offline and online, specialist and generalist, as well as cold-start and warm-start scenarios. For example, the online alignment experiments showed that CUT can improve LLMs in an iterative fashion with up-to-date, model-specific judgments, akin to how humans progressively refine their behaviors through ongoing feedback from their peers over time. Our analysis comparing rewards and judgments suggested that aligning LLMs with judgments is a promising research area.
Appendix A Appendix
A.1 Alignment Templates
Figure 5 shows the templates when we apply CUT to align LLMs. Figure 6 shows the inference template, which does not necessitate judgments.
A.2 Prompt for Judgment Annotation
Figure 7 illustrates the prompt employed to request GPT-4’s assistance in annotating judgments. We consider the judgment that begins with the keyword ”Perfect.” to be a positive judgment; otherwise, it is deemed a negative judgment. GPT-4 demonstrates proficiency in fulfilling this requirement. Figure 8 shows the template used for training AI judges.
Figure 5: The template used for aligning LLMs through CUT.
Figure 6: The inference template used for LLaMA2 and LLaMA2-chat respectively.
Figure 7: The prompt for asking GPT4 in annotating judgment.
Figure 8: The template used for training AI judges.

### v2

@@ 5.2.1 Iterative Alignment
Setup. As mentioned in § 3.1, the online alignment process can be conducted iteratively, akin to how humans continuously refine their behaviors through ongoing feedback. Specifically, we apply the following three steps repeatedly:
  • Step 1: Collect a set of instructions $\bm{x}$, and obtain the responses $\bm{y}$ from the target model.
@@
where the target LLM is LLaMA2-chat. In each iteration, we sample 1000 distinct instructions from Stanford Alpaca (Taori et al., 2023). We ask GPT4 for drafting judgments, which has been proven to produce high-quality annotations (Cui et al., 2023). Annotation details are elaborated in Appendix A.2. Note that most responses from LLaMA2-chat receive positive judgments, resulting in a large proportion of Align-P examples. We found downsampling Align-P examples is beneficial to the online alignment (see Appendix A.4). We evaluate models on ARC, HellaSwag, MMLU, TruthfulQA, and AlpacaEval.
Figure 3: The results of online alignment with different AI judges.
Results. Table 6 shows the results of online iterative alignment. In the first iteration, online alignment exhibits superior performance over offline alignment on both TruthfulQA and AlpacaEval. This observation implies that model-specific judgments are more effective for alignment. More importantly, the alignment continues to improve with more iterations, where the performance rises from 81.09 to 91.68 on AlpacaEval after four iterations. However, the performance improvement ceases at the fifth iteration. We speculate two possible explanations for this occurrence: (1) the judgments provided by GPT-4 contain certain inaccuracies, making them insufficient to effectively align a strong target model like our CUT 4+. (2) The target model may exhibit a knowledge deficiency in specific domains, such as mathematics and science, which cannot be adequately addressed through judgments. We also provide a case study in Appendix A.5.
5.2.2 Training A Judgment Model
In the previous experiments, we show that CUT is effective in aligning LLMs with judgments annotated by humans or GPT4. However, both annotations can be expensive or infeasible. Therefore, we investigate the possibilities of developing an AI judge based on the current open-source LLMs.
Setup. we train AI judges with different amounts of judgment data $\{3000,5000\}$ collected in § 5.2.1. Then, we sample 1000 new instructions from Stanford Alpaca, obtain the corresponding responses from the target model (i.e., LLaMA2-chat), and label judgments with our AI judges. These new judgment triplets are used to align the target model.
Results. Figure 3 shows that AI judge-5000, trained with 5000 judgment data, is beneficial for aligning the target LLM, which leads to improvements of 3.02 and 4.17 points compared to LLaMA2-chat on TruthfulQA and AlpacaEval respectively. In contrast, AI Judge-3000, using a smaller training dataset, shows limited effectiveness. The comparison suggests that training a capable AI judge necessitates a moderate number of high-quality training instances. As a result, it is feasible to train AI judges to align the LLM. However, the quality of the AI judge remains a crucial factor in determining the success of this endeavor.
Figure 4: Comparison between reward-based DPO and judgment-based CUT.
5.3 Judgment vs. Reward
Our work primarily focuses on aligning LLMs with judgments, whereas most prior research explores rewards. In this section, we aim to provide a direct comparison between these two paradigms. However, note that it is hard to conduct a fair comparison due to the distinct data formats and the potential variation in data quality.
Setup. We compare judgment-based CUT with the state-of-the-art reward-based DPO (Rafailov et al., 2023). To maximize fairness, we leverage UltraFeedback (Cui et al., 2023), which contains both reward and judgment annotations produced by GPT4. Our preliminary experiments show that CUT is not good using the original judgments in UltraFeedback. We find that the reason is that the judgments in UltraFeedback tend to commend the strengths of given responses. This type of judgment is unsuitable for our CUT, as we primarily use judgments for inappropriate token detection. Therefore, we re-collect judgments on the same instruction-response pairs from GPT4 using our prompt (Appendix A.2). Due to budget constraints, we randomly sample up to 3000 instructions (with 4 responses each, totaling 12,000 pairs) for annotation. The implementation details are as follows:
  • DPO: For each of the above instructions, we formulate preference data by enumerating all possible pairs of responses from the given four, excluding pairs with the same reward value.
  • CUT-UF: We fine-tune the base model on the above instruction-response pairs and their original judgments from UltraFeedback using CUT.
  • CUT: We use the same instruction-response pairs as CUT-UF but with our re-annotated judgments.
Results. Figure 4 (a) shows the effect of three alignment methods using 1000 instructions as the alignment data. We can observe that CUT consistently improves over CUT-UF on all five tasks for two base models, which verifies our assumption that CUT is more effective when using critics as the judgment. Notably, CUT surpasses DPO by a large margin of 37.54 and 23.04 points on AlpacaEval for two base models, respectively. This shows that CUT is more effective in aligning LLMs with limited alignment data (i.e., 1000 instructions). Figure 4 (b) depicts the trends when adding more data for CUT and DPO alignment. The performance of CUT on these tasks is generally better or comparable to that of DPO and demonstrates a positive correlation with the size of the training data provided. The above observations suggest that judgments hold greater potential than rewards in aligning LLMs. CUT is slightly worse than DPO on ARC, and HellaSwag. We hypothesize that the performance discrepancy is partly caused by the evaluation protocols: the four tasks are ranking-based. As suggested Bansal et al. (2023), methods such as DPO, which leverage ranking data in the alignment possess inherent advantages in ranking-based tasks. We also provide a case study in Appendix A.6.
6 Conclusion
We systematically explored the alignment of LLMs through the lens of judgments. We investigated three potential methods that can be adapted for aligning LLMs with judgments but found them unable to fully capitalize on judgments. We proposed a novel framework CUT, that enables direct and explicit learning from judgments and facilitates fine-grained inappropriate content detection and correction. Extensive evaluations demonstrated the effectiveness of our CUT in various settings, including offline and online, specialist and generalist, as well as cold-start and warm-start scenarios. For example, the online alignment experiments showed that CUT can iteratively improve LLMs with up-to-date model-specific judgments, akin to how humans progressively refine their behaviors through ongoing feedback. Our analysis comparing rewards and judgments suggested that aligning LLMs with judgments is a promising research area.
Limitations
Quality of Judgment Models
  • In an attempt to augment the training size, we incorporated the 1317 judgment data from Shepherd for training the AI judge. However, after including Shepherd, the AI judge’s performance deteriorated, resulting in more illogical judgments such as "The original answer 100 is incorrect. The correct answer should be 100." We hypothesize that reasoning and math tasks from Shepherd are too complex for a 13b model to comprehend. Consequently, larger language models may be required to achieve better judgment generation quality, a notion supported by Saunders et al. (2022).
Size of Alignment Data
Due to budgetary constraints, our research currently involves experiments utilizing several thousands of judgment data. In future research endeavors, we would like to investigate the scaling law with an expanded volume of judgment data.
Appendix A Appendix
A.1 Implementations
We train our models using LoRA (Hu et al., 2022) and follow the best configurations suggested by Platypus (Lee et al., 2023). The tradeoff hyperparameter $\lambda$ is selected from $\{1.1,1.2\}$ and the unlikelihood weight $\alpha$ and $\gamma$ is selected from $\{0.25,0.5,1\}$ and $\{0.25,0.5,1,2\}$, respectively. We adopt the Alpaca template (Taori et al., 2023) for fine-tuning and inference. Figure 5 shows the templates when we apply CUT to align LLMs. Figure 6 shows the inference template, which does not necessitate judgments.
A.2 Prompt for Judgment Annotation
Figure 8 illustrates the prompt employed to request GPT-4’s assistance in annotating judgments. We consider the judgment that begins with the keyword "Perfect." to be a positive judgment; otherwise, it is deemed a negative judgment. GPT-4 demonstrates proficiency in fulfilling this requirement. Figure 9 shows the template used for training AI judges.
Figure 5: The template used for aligning LLMs through CUT.
Figure 6: The inference template.
@@ A.2 Prompt for Judgment Annotation
Figure 8: The prompt for asking GPT4 in annotating judgment.
Figure 9: The template used for training AI judges.
A.3 Offline Alignment Tasks
  • General Instruction-following: We train models on the Shepherd dataset (Wang et al., 2023a), which consists of judgment data on diverse NLP tasks such as math word problems and commonsense reasoning. There are 1317 examples in total. For evaluation, we report model performance on four ranking-based and one generation-based LLM benchmarks, where ranking-based evaluation tests an LLM’s ability to select the best response from a set of candidate responses, while generation-based evaluation assesses an LLM’s ability to generate high-quality responses. Following the Open LLM Leaderboard (Gao et al., 2021), the ranking-based benchmarks are 25-shot ARC (Clark et al., 2018), 10-shot HellaSwag (Zellers et al., 2019), 5-shot MMLU (Hendrycks et al., 2021), and 0-shot TruthfulQA (Lin et al., 2022) from the Open LLM Leaderboard (Gao et al., 2021). The generation-based benchmark is AlpacaEval [^2] Following conventions, GPT4 is utilized to judge the winning rate of the responses generated by our models against those produced by DaVinci003..
A.4 Downsampling Align-P
Mixing training data from different categories can substantially affect the performance of trained models (Xu et al., 2023a, b, c). As LLaMA2-chat has already undergone extensive alignment training, its responses to the Stanford Alpaca instructions are generally of high quality. In fact, 713 out of 1000 responses generated by LLaMA2-chat receive positive judgments, resulting in a substantial proportion of Align-P examples. To investigate the effect of the proportion of Align-P examples, we undertake a downsampling process for these examples. The performance of various downsampling ratios is illustrated in Figure 7. Our findings indicate that maintaining a moderate percentage of Align-P examples is crucial. We conjecture that preserving a certain number of Align-P examples allows the model to sustain its capacity to generate satisfactory responses, while too many Align-P examples may lead to overfitting, thereby disrupting the alignment process. In subsequent experiments, we keep a ratio of 0.25.
A.5 Case Study: Online Alignment
Instruction: What color is the sky?
LLaMA2-chat: The color of the sky is blue.
CUT 1+: The color of the sky is blue.
CUT 2+: The color of the sky is blue.
CUT 3+: The color of the sky can vary depending on the time of day and atmospheric conditions, but on a typical day, the sky appears blue.
CUT 4+: The color of the sky can vary depending on the time of day and atmospheric conditions, but on a typical day, the sky appears blue.
@@ A.5 Case Study: Online Alignment
LLaMA2-chat: The output of the given LATEX code is a document classified as an "article" with the file name "First document.tex". The document contains only one line of text "First document. This is a simple example, with no extra parameters or packages included.
@@
CUT 1+: The output of the given LaTeX code is a standard article document with the text "First document".
@@
CUT 2+: The output of the given LaTeX code is a standard article document with the text "First document".
@@
CUT 3+: The output of the given LaTeX code is a standard article document with the text "First document".
@@
CUT 4+: Sure! Here’s the output of the given LaTeX code: First document. This is a simple example, with no extra parameters or packages included.
@@
Instruction: Write a good Twitter Bio. Try to cover all the provided main points in your generated bio. Name: Jane Main points: - Architect - Interested in Tech Written in first person Use casual tone
@@
LLaMA2-chat: Hey there! I’m Jane, a passionate architect with a love for all things tech. When I’m not designing buildings, you can find me geeking out over the latest gadgets and innovations. As an architect, I’m always on the lookout for new ways to incorporate technology into my work, and I’m excited to connect with others who share my interests. Let’s build something amazing together! #tech #architecture #innovation
@@
CUT 1+: Hey there, I’m Jane! I’m an architect with a passion for all things tech. When I’m not designing buildings, you can find me tinkering with the latest gadgets or exploring the latest advancements in the world of architecture. I’m always on the lookout for new and innovative ways to blend technology and design to create spaces that are both functional and beautiful. Follow me for a mix of architecture, tech, and design inspiration!
@@
CUT 2+: Hey there, I’m Jane! I’m an architect with a passion for all things tech. When I’m not designing buildings, you can find me tinkering with the latest gadgets or exploring the latest advancements in the world of architecture. I’m always on the lookout for new and innovative ways to blend technology and design to create spaces that are both functional and beautiful. Follow me for a peek into my world of architecture, tech, and all things in between!
@@
CUT 3+: Hey there, I’m Jane! I’m an architect with a passion for all things tech. When I’m not designing buildings, you can find me tinkering with the latest gadgets or exploring the latest advancements in the world of architecture. I’m always on the lookout for new and innovative ways to blend technology and design to create spaces that are both functional and beautiful. Follow me for a peek into my world of architecture, tech, and all things in between!
@@
CUT 4+: Hey there, I’m Jane! I’m an architect with a passion for all things tech. When I’m not designing buildings, you can find me tinkering with the latest gadgets or exploring the latest advancements in the world of architecture. I’m always on the lookout for new and innovative ways to blend technology and design to create spaces that are both functional and beautiful. Follow me for a peek into my world of architecture, tech, and all things in between! #architect #tech #innovation #design
@@
A.6 Case Study: CUT v.s. DPO
Instruction: How did US states get their names?
@@
DPO: Good question! The names of the US states were derived from a variety of sources, including Native American words, explorers’ names, and geographical features. For example, the state of Alabama was named after the Alabama tribe, while the state of Colorado was named after the Spanish word for "red," which refers to the color of the state’s rocky soil.
@@
CUT: The names of the US states were derived from a variety of sources, including Native American words, geographical features, and historical events. Here are some examples of how some of the states got their names: 1. Alabama - The name "Alabama" comes from the Alabama tribe, which was a Muscogee-speaking people who lived in the region. 2. Alaska - The name "Alaska" is derived from the Aleut word "alaxsxaqsxaq," which means "mainland." 49. Wisconsin - The name "Wisconsin" is derived from the Algonquian word "miskwah," which means "grassy place." 50. Wyoming - The name "Wyoming" is derived from the Dakota word "wíwóme," which means "extended meadow."
@@
Instruction: I’m trying to teach myself to have nicer handwriting. Can you help?
@@
DPO: Of course! I’d be happy to help you improve your handwriting. What specific areas are you looking to work on? Do you want to focus on letter formation, spacing, or overall legibility?
CUT: Of course! Improving your handwriting can be a fun and rewarding process. Here are some tips to help you achieve nicer handwriting: 1. Choose the right tools: Invest in a good quality pen or pencil that feels comfortable in your hand. Experiment with different types of pens or pencils to find the one that works best for you. 2. Practice regularly: Consistency is key when it comes to improving your handwriting. Set aside some time each day to practice writing. Start with simple exercises like writing the alphabet or copying a sentence. 10. Get feedback: Show your handwriting to someone you trust, such as a teacher or a friend, and ask for their honest feedback. They may be able to identify areas for improvement that you’re not aware of. Remember, improving your handwriting takes time and practice. Be patient and persistent, and you’ll see improvement over time. Good luck!
Instruction: Write "Test"
DPO: Sure! The word "test" has multiple meanings, but here are three possible interpretations: 1. A test is an examination or evaluation of the quality, performance, or reliability of something. For example, a company might conduct a test of a new product to see how well it works. 2. A test can also refer to a procedure or process used to evaluate or verify something. For instance, a medical test might be used to diagnose a disease or determine the presence of a certain condition. 3. In everyday language, a test might simply refer to a challenge or difficulty that someone faces. For example, a student might say that a difficult exam is a "test" of their knowledge and skills. Is there anything else you’d like to know about the word "test"?
CUT: Test